# A Caveat on Metrizing Convergence in Distribution on Hilbert Spaces

### v2

@@ 1 Introduction
In [2], the second and third authors of the present paper introduced a probabilistic distance, noted $d_{2}(X,Y)$ and formally identified in Definition 2 and Remark 1 below, between the distributions of two random elements $X,Y$ taking values in a real separable Hilbert space $K$. As we will see, such a distance is obtained by considering the supremum of the quantity $|\mathbb{E}[f(X)]-\mathbb{E}[f(Y)]|$, where $f$ runs over a certain class of twice Fréchet differentiable test functions $f$, such that the second derivative $D^{2}f$ induces a Hilbert-Schmidt operator on $K$ (see Definition 1 for a precise formulation). The distance $d_{2}$ is claimed in [2, Lemma 3.1] to metrize convergence in distribution on $K$.
The use of the distance $d_{2}$ allowed the authors of [2] to develop an infinite-dimensional version of the so-called Malliavin-Stein method (see e.g. [19]), including Fourth Moment and Breuer-Major type quantitative CLTs. Further uses of the distance $d_{2}$, and of the techniques initiated in [2], have appeared in subsequent years (and are continuing to appear at the time of writing), spanning fields as diverse as stochastic analysis on configuration spaces, the geometry of random fields, and the large-scale study of randomly initialized neural networks — see [3, 4, 6, 7, 14, 15, 21] [^1] We stress that $d_{2}$ is only needed for [15, formula (3.27)]; all other results in [15] are independent of both $d_{2}$ and formula (3.27) therein. .
The purpose of this paper is to provide a formal counterexample to [2, Lemma 3.1], by showing that if ${\rm dim}\,K=+\infty$, the distance $d_{2}$ cannot be employed to deduce convergence in distribution on $K$.  [^2] The proof of [2, Lemma 3.1] is based on an erroneous interpretation of a result in [8], where the considered distance is not the distance $d_{2}$, but rather the distance $\rho_{\infty}$ defined below.
@@
As a consequence of Theorem 1 (note that $\rho_{2}=d_{2}$, see Remark 1), the proofs of the functional CLTs deduced in the references [3, 4, 6, 7, 14, 15] through the use of the distance $d_{2}$ are not valid. The proof of Theorem 1 (as given in Section 3.2 below) is by contradiction, and exploits the fact that the metrization property of $\rho_{p}$, for any $p\in[1,\infty)$, would negate some classical criteria for the convergence in distribution of Hilbert space-valued Gaussian random elements, see [1, Chapter 3].
@@ 2.1 Linear mappings and bilinear forms
In this section, we recall some basic notions about linear mappings and bilinear forms on Banach and Hilbert spaces. Our main reference is [18, Chapters IV and XIII]; see also [11, 12, 22, 5]
@@
so that $(L(E,F;G),\|\cdot\|)$ is also a Banach space (since $G$ is Banach). It is a classical result (see e.g. [18, p. 67]) that $L(E,F;G)\simeq L(E,L(F,G))$ where $\simeq$’ indicates the existence of a Banach space isomorphism (that is, a norm-preserving, continuous linear mapping whose inverse is also continuous and linear). When $E=F$, we will say that $\varphi\in L(E,E;G)$ is symmetric if $\varphi(x,y)=\varphi(y,x)$, for all $x,y\in E$.
@@
where $\{e_{i}\}$ is an arbitrary orthonormal basis of $K$ (the definition of $\langle\cdot,\cdot\cdot\rangle_{HS}$ is independent of the choice of the basis); we will henceforth use the symbol $\mathcal{HS}(K)$ to denote such a Hilbert space. We will also write $K\otimes K$ and $K\odot K$, respectively, to indicate the tensor product and the symmetric tensor product of $K$ with itself (see e.g. [17, Appendix E] for a concise and probability-oriented discussion of tensor products).
@@ 2.2 Differentiable mappings
Then, $f\in C^{2}(K,\mathbb{R})$. Moreover, the second derivative of $f$ is compact if and only if $\psi$ is constant.
@@ 3.1 Probabilistic distances on Hilbert spaces
The distance $\rho_{\infty}$ is an element (up to some minor variations) of the class of discrepancies considered in [8, 16] (see also [13]). It is a well-known fact (see e.g. [16, Theorem 2.4] or [8, Theorem 7]) that $\rho_{\infty}$ metrizes convergence in distribution on every separable real Hilbert space $K$. As previously discussed, Theorem 1 — whose proof is the subject of the next section — establishes that for $p\in[1,\infty)$, the distance $\rho_{p}$ metrizes convergence in distribution only in the trivial finite-dimensional case, that is, when $K$ has finite dimension and the choice of norm is thus irrelevant.
@@ 3.2 Proof of Theorem 1
We will also need the following estimate, which is a generalization of [15, Proposition 5.12].
@@
Let $h\in\mathcal{F}_{p}$ and define $U_{t}=\sqrt{t}X_{1}+\sqrt{1-t}X_{2}$ for $t\in[0,1]$. Proceeding as in the proof of [15, Proposition 5.12], we obtain
$\mathbb{E}\left[X_{1}\right]-\mathbb{E}\left[X_{2}\right]=\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\left\langle D^{2}h(U_{t}),S_{1}-S_{2}\right\rangle_{HS}\right]dt$
@@
$\left|\mathbb{E}\left[X_{1}\right]-\mathbb{E}\left[X_{2}\right]\right|$ $\leq\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\operatorname{Tr}\left|\left(D^{2}h(U_{t})\right)^{\ast}\left(S_{1}-S_{2}\right)\right|\right]dt$
@@ 4 Proof of Lemma 2
Using the standard chain rule (as stated e.g. in [18, p. 337]), one has that
@@
so that $f\in C^{1}(K,\mathbb{R})$. Applying the product rule (see [18, p. 336]) yields also that, for all $x,h,k\in K$,

### v3

@@ 1 Introduction
In [3], the second and third authors of the present paper introduced a probabilistic distance, noted $d_{2}(X,Y)$ and formally identified in Definition 2 and Remark 1 below, between the distributions of two random elements $X,Y$ taking values in a real separable Hilbert space $K$. As we will see, such a distance is obtained by considering the supremum of the quantity $|\mathbb{E}[f(X)]-\mathbb{E}[f(Y)]|$, where $f$ runs over a certain class of twice Fréchet differentiable test functions $f$, such that the second derivative $D^{2}f$ induces a Hilbert-Schmidt operator on $K$ (see Definition 1 for a precise formulation). The distance $d_{2}$ is claimed in [3, Lemma 3.1] to metrize convergence in distribution on $K$.
The use of the distance $d_{2}$ allowed the authors of [3] to develop an infinite-dimensional version of the so-called Malliavin-Stein method (see e.g. [18]), including Fourth Moment and Breuer-Major type quantitative CLTs. Further uses of the distance $d_{2}$, and of the techniques initiated in [3], have appeared in subsequent years (and are continuing to appear at the time of writing), spanning fields as diverse as stochastic analysis on configuration spaces, the geometry of random fields, and the large-scale study of randomly initialized neural networks — see [2, 4, 6, 7, düker2024breuermajortheoremshilbertspacevalued, 14, 20] [^1] We stress that $d_{2}$ is only needed for [14, formula (3.27)]; all other results in [14] are independent of both $d_{2}$ and formula (3.27) therein. .
The purpose of this paper is to provide a formal counterexample to [3, Lemma 3.1], by showing that if ${\rm dim}\,K=+\infty$, the distance $d_{2}$ cannot be employed to deduce convergence in distribution on $K$.  [^2] The proof of [3, Lemma 3.1] is based on an erroneous interpretation of a result in [8], where the considered distance is not the distance $d_{2}$, but rather the distance $\rho_{\infty}$ defined below.
@@
As a consequence of Theorem 1 (note that $\rho_{2}=d_{2}$, see Remark 1), the proofs of the functional CLTs deduced in the references [2, 4, 6, 7, düker2024breuermajortheoremshilbertspacevalued, 14] through the use of the distance $d_{2}$ are not valid. The proof of Theorem 1 (as given in Section 3.2 below) is by contradiction, and exploits the fact that the metrization property of $\rho_{p}$, for any $p\in[1,\infty)$, would negate some classical criteria for the convergence in distribution of Hilbert space-valued Gaussian random elements, see [1, Chapter 3].
@@ 2.1 Linear mappings and bilinear forms
In this section, we recall some basic notions about linear mappings and bilinear forms on Banach and Hilbert spaces. Our main reference is [17, Chapters IV and XIII]; see also [11, 12, 21, 5]
@@
so that $(L(E,F;G),\|\cdot\|)$ is also a Banach space (since $G$ is Banach). It is a classical result (see e.g. [17, p. 67]) that $L(E,F;G)\simeq L(E,L(F,G))$ where $\simeq$’ indicates the existence of a Banach space isomorphism (that is, a norm-preserving, continuous linear mapping whose inverse is also continuous and linear). When $E=F$, we will say that $\varphi\in L(E,E;G)$ is symmetric if $\varphi(x,y)=\varphi(y,x)$, for all $x,y\in E$.
@@
where $\{e_{i}\}$ is an arbitrary orthonormal basis of $K$ (the definition of $\langle\cdot,\cdot\cdot\rangle_{HS}$ is independent of the choice of the basis); we will henceforth use the symbol $\mathcal{HS}(K)$ to denote such a Hilbert space. We will also write $K\otimes K$ and $K\odot K$, respectively, to indicate the tensor product and the symmetric tensor product of $K$ with itself (see e.g. [16, Appendix E] for a concise and probability-oriented discussion of tensor products).
@@ 2.2 Differentiable mappings
Then, $f\in C^{2}(K,\mathbb{R})$. Moreover, for each $x\in K$ the second Fréchet derivative $D^{2}f(x)$ (a bounded bilinear form on $K\times K$) can be identified via the Riesz isomorphism with a bounded self-adjoint operator on $K$. In this sense, $D^{2}f(x)$ is a compact operator on $K$ for every $x\in K$ if and only if $\psi$ is constant.
@@ 3.1 Probabilistic distances on Hilbert spaces
The distance $\rho_{\infty}$ is an element (up to some minor variations) of the class of discrepancies considered in [8, 15] (see also [13]). It is a well-known fact (see e.g. [15, Theorem 2.4] or [8, Theorem 7]) that $\rho_{\infty}$ metrizes convergence in distribution on every separable real Hilbert space $K$. As previously discussed, Theorem 1 — whose proof is the subject of the next section — establishes that for $p\in[1,\infty)$, the distance $\rho_{p}$ metrizes convergence in distribution only in the trivial finite-dimensional case, that is, when $K$ has finite dimension and the choice of norm is thus irrelevant.
@@ 3.2 Proof of Theorem 1
We will also need the following estimate, which is a generalization of [14, Proposition 5.13].
@@
Let $h\in\mathcal{F}_{p}$ and define $U_{t}=\sqrt{t}X_{1}+\sqrt{1-t}X_{2}$ for $t\in[0,1]$. Proceeding as in the proof of [14, Proposition 5.13], we obtain
$\mathbb{E}\left[h\left(X_{1}\right)\right]-\mathbb{E}\left[h\left(X_{2}\right)\right]=\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\left\langle D^{2}h(U_{t}),S_{1}-S_{2}\right\rangle_{HS}\right]dt$
@@
$\left|\mathbb{E}\left[h\left(X_{1}\right)\right]-\mathbb{E}\left[h\left(X_{2}\right)\right]\right|$ $\leq\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\operatorname{Tr}\left|\left(D^{2}h(U_{t})\right)^{\ast}\left(S_{1}-S_{2}\right)\right|\right]dt$
@@ 4 Proof of Lemma 2
Using the standard chain rule (as stated e.g. in [17, p. 337]), one has that
@@
so that $f\in C^{1}(K,\mathbb{R})$. Applying the product rule (see [17, p. 336]) yields also that, for all $x,h,k\in K$,